What will come in place of question mark (?) in the given number series ? 14 7 7 14 55 448

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiii Arun 
The ans would be 7 
14 * 0.5 = 7
7* 1 = 7
7* 2 = 14
14*4 = 56
56* 8 = 448 
so , 7 is correct ans.